OpenStudy (anonymous):

HELP PLEASE If f(x) = x − 4 and g(x) = x − 2, find f[g(−6)].

OpenStudy (anonymous):

Plug -6 into g(x) first, than plug the anwser in for x in f(x)

OpenStudy (anonymous):

??

OpenStudy (ledah):

Can't get any simpler than his/her explanation. :c

OpenStudy (anonymous):

DIDNT SAY IT CULD LOL

OpenStudy (anonymous):

g(-6) = - 6 - 2 = ? plug ? into x in the equation f(x) = x - 4 to get f(?) = ? - 4 than add like terms. lol

OpenStudy (ledah):

Well the double question marks make it seem like you just don't get it

OpenStudy (anonymous):

U WROTE CANT GET ANY SIMPLER BFORE I DID THE TWO ??

OpenStudy (ledah):

after actually.

OpenStudy (anonymous):

THEMUSICIAN I GOT THE ANSWER ITS 9 RIGHT/? BTW LEDAH I GOT OF URE CONVO U GET OFF MY QUESTION PLZ!!!

OpenStudy (ledah):

was actually only here to help. seeing as I've done nothing wrong. and that wasn't my conversation by the way. that was just the chat box for EVERYONE to use.

OpenStudy (anonymous):

No. -6 -2 = -8 plug that into f(x) = x - 4 it becomes -8 -4 = ?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

OK LEDDAH IF THATS THE CHAT BOX FOR EVERYONE TO USE HOW COME U TOLD ME TO GET OFF OR BACKOFF OR STOP OR W.E U SAID

OpenStudy (anonymous):

No. a Negative + a negative is a negative. -8 -4 = -8 + (-4) = - 16. Do you get what I did?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

-12**

OpenStudy (anonymous):

Sorry typo. It is -12*

OpenStudy (anonymous):

OOPS I WROTE THE QUESTION WRONG THE QUESTION IS f f(x) = x − 4 and g(x) =1/2 x − 2, find f[g(−6)]

OpenStudy (ledah):

why are you insisting on using caps. and who said I was talking to you about backing off. I only said stop copying people.. get your accusations right.

OpenStudy (anonymous):

SHUT UP AND BYE I HAVE BETTER STUFF TO DO THEN THINK ABOUT MY GRAMMER NOW IF U EXCUSE ME ME AND THEMUSICIAN ARE DOING WORK

OpenStudy (anonymous):

Oooh. Okay. So (1/2)(-6) - 2 = -3 -2 = -5 plug that into x - 4 which becomes -5 -4 =?

OpenStudy (anonymous):

Now do you see what I did?

OpenStudy (anonymous):

YES :)

OpenStudy (anonymous):

:] Okay. So what do you think the anwser is?

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

Yes :D

OpenStudy (anonymous):

THANK U

OpenStudy (anonymous):

No problem :]